OpenStudy (anonymous):

a 68 degrees F, a cricket chirps 24 times per min. at 40 degrees F, the same cricket chirps 86 times per min. express the number of chirps (y) as a linear function of the Fahrenheit temperature (x)

OpenStudy (amistre64):

we have two "coordinates" we can use to find a "slope". 24,68 and 86,40 68-40 ------ = slope 24-86 28/-62 = -14/31 better flip that to get y chirps to degrees y = -31/24x + b 24 = -31/14(68) + b b = 24 + 31(68)/14 y = -31/14x + 174.57 if I did it right :)

OpenStudy (anonymous):

thank you